OpenStudy (anonymous):

within a group of n people, the number of possible handshakes,N, is given by N=1/2(n^2-n). How many handshakes are possible at a meeting if 30 people are present?

OpenStudy (anonymous):

can't be

OpenStudy (anonymous):

i'm lost

OpenStudy (anonymous):

somewhere along the lines of, 420,870,465,435

hero (hero):

N = (30^2-30)/2 = (900 - 30)/2 = 870/2 = 435 Handshakes

OpenStudy (anonymous):

thank you sir!

hero (hero):

Sorry, read it wrong at first

OpenStudy (anonymous):

no problem

OpenStudy (zarkon):

I prefer to write it as \[_{30}C_{2} \text{ or }{30\choose 2}\]

hero (hero):

Yeah...but

OpenStudy (zarkon):

my calculator has a built in function ncr(30,2)=435 quick and easy

hero (hero):

Isn't that nice?

hero (hero):

I don't know where you got the 2 though

OpenStudy (zarkon):

there are 30 people...pick 2 at a time to shake hands where order does not matter.